# Collusion-Resilience in Transaction Fee Mechanism Design

### v2

@@ 7.1 Allowing the Globally Optimal Strategy to Coordinate
The mechanism is parametrized with some reserve price $r>0$ and block size $k$. Any bid that is at least $r$ is considered eligible. Randomly select up to $k$ bids to include in the block. All included bids are confirmed and they each pay $r$. All payment is burnt and the miner receives nothing. It is not hard to see that the above mechanism satisfies UIC and MIC. It also satisfies OCA-proofness with the aforementioned relaxation since the global coalition can just have the users with the top $k$ values bid more than $r$, while everyone else bids 00.
@@ 7.2 Allowing the Globally Optimal Strategy to Output Multiple Bids
    • Case 3: Else, if the number of bids is odd, henceforth denoted $12n+$ 1, check if there are $n$ bids at $r$ and $n$ bids in the range of $(0,1]$, and one last bid denoted $r^{\prime}$. If $r^{\prime}>r$, then choose $r^{\prime}$ to include in the block, and among the remaining $n$ bids at $r$, randomly choose $\min(k-1,n)$ of them to include in the block; else if $r^{\prime}\leq r$, randomly choose $\min(k,n)$ bids among all the $r$-bids to include in the block.
@@
    Second, before user $i$ submits its bid, the bid vector does not contain exactly $n$ bids at $r$ and $n$ bids lie in the range $(0,1]$ for some $n$. Suppose there is no bid strictly larger than $r$ in the bid vector before user $i$ submits its bid. Then, if user $i$ bids truthfully, its bid will be the only bid larger than $r$, and the inclusion rule goes to Case 4. Since $k\geq 1$, user $i$’s bid is guaranteed to be included and confirmed, and thus maximizes the utility. On the other hand, suppose there are $t\geq 1$ bids strictly larger than $r$ in the bid vector before user $i$ submits its bid. Then, the inclusion can never go to Case 1 or Case 2. If user $i$ bids truthfully, the inclusion rule must go to Case 4, and user $i$’s bid is included with probability $\min(k,t+1)/(t+1)$. However, if the strategic bidding leads to Case 3, it must be $t=1$. Then, user $i$’s inclusion probability has dropped from $1/2$ to 00 if $k=1$, or from $1$ to at most 1 if $k\geq 2$. On the other hand, if the strategic bidding leads to Case 4, any untruthful bidding $>r$ leads to the same inclusion probability and the same payment, so the utility does not change, while any untruthful bidding $\leq r$ leads to the zero inclusion probability which leads to zero utility. Finally, notice that injecting fake bids never helps in Case 4. Thus, in all cases, the utility is better than bidding truthfully.

### v3

@@ 7.1 Allowing the Globally Optimal Strategy to Coordinate
The mechanism is parametrized with some reserve price $r>0$ and block size $k$. Any bid that is at least $r$ is considered eligible. Randomly select up to $k$ bids to include in the block. All included bids are confirmed and they each pay $r$. All payment is burnt and the miner receives nothing. It is not hard to see that the above mechanism satisfies UIC and MIC. It also satisfies OCA-proofness with the aforementioned relaxation since the global coalition can just have the users with the top $k$ values bid more than $r$, while everyone else bids 0.
@@ 7.2 Allowing the Globally Optimal Strategy to Output Multiple Bids
    • Case 3: Else, if the number of bids is odd, henceforth denoted $2n+1$, check if there are $n$ bids at $r$ and $n$ bids in the range of $(0,1]$, and one last bid denoted $r^{\prime}$. If $r^{\prime}>r$, then choose $r^{\prime}$ to include in the block, and among the remaining $n$ bids at $r$, randomly choose $\min(k-1,n)$ of them to include in the block; else if $r^{\prime}\leq r$, randomly choose $\min(k,n)$ bids among all the $r$-bids to include in the block.
@@
    Second, before user $i$ submits its bid, the bid vector does not contain exactly $n$ bids at $r$ and $n$ bids lie in the range $(0,1]$ for some $n$. Suppose there is no bid strictly larger than $r$ in the bid vector before user $i$ submits its bid. Then, if user $i$ bids truthfully, its bid will be the only bid larger than $r$, and the inclusion rule goes to Case 4. Since $k\geq 1$, user $i$’s bid is guaranteed to be included and confirmed, and thus maximizes the utility. On the other hand, suppose there are $t\geq 1$ bids strictly larger than $r$ in the bid vector before user $i$ submits its bid. Then, the inclusion can never go to Case 1 or Case 2. If user $i$ bids truthfully, the inclusion rule must go to Case 4, and user $i$’s bid is included with probability $\min(k,t+1)/(t+1)$. However, if the strategic bidding leads to Case 3, it must be $t=1$. Then, user $i$’s inclusion probability has dropped from $1/2$ to 0 if $k=1$, or from $1$ to at most 1 if $k\geq 2$. On the other hand, if the strategic bidding leads to Case 4, any untruthful bidding >r>r leads to the same inclusion probability and the same payment, so the utility does not change, while any untruthful bidding $\leq r$ leads to the zero inclusion probability which leads to zero utility. Finally, notice that injecting fake bids never helps in Case 4. Thus, in all cases, the utility is better than bidding truthfully.